OpenStudy (yumeybitaa):

Find -a2 - 3b3 + c2 + 2b3 - c2 if a = 3, b = 2, and c = -3. -67 -17 -31 1

OpenStudy (nirmalnema):

-(3)^2 - 3(2)^3 + (-3)^2 + 2(2)^3 - (-3)2 = -9-24+9-9+16 =-17